OpenStudy (koikkara):

What is 10 to the power of 12 called?

OpenStudy (amistre64):

10 to the 12th

OpenStudy (amistre64):

...or are you asking what a million million is refered to as?

OpenStudy (dmezzullo):

Most likely the 1st one

OpenStudy (koikkara):

hmm...no 10^9=Nano....similarily, 10^12=?

terenzreignz (terenzreignz):

Nobody calls it trillion?

OpenStudy (koikkara):

exactly, then what to call...?

OpenStudy (anonymous):

I'm assuming you meant 10^(-9) = Nano So I think you're talking about 10^(-12), which would be called pico

OpenStudy (koikkara):

10 raised by the ten power is 10000000000, also known as ten million....then...??

OpenStudy (anonymous):

10^12 = 1000000000000, also known as a trillion

OpenStudy (koikkara):

sure....

terenzreignz (terenzreignz):

http://en.wikipedia.org/wiki/Long_and_short_scales

OpenStudy (koikkara):

thanks tenzz....for the link....

OpenStudy (koikkara):

@Meepi @terenzreignz @amistre64 @dmezzullo ......i applaud u guys, for your valuable help....Thanks guys...!!

OpenStudy (anonymous):

in computing and engineering you will be using Tera